OpenStudy (anonymous):

Find the remainder when x51+51 is divided by (x+1).

OpenStudy (jack1):

heya

OpenStudy (anonymous):

do u know the solution for this problem......jack

OpenStudy (anonymous):

if F(x) is divided by x+a the remainder is given by F(a)

OpenStudy (jack1):

x51+51 is divided by (x+1) is that \( \Large \frac {51x+51}{ (x+1)}\)? or is it\( \Large \frac {x^{51}+51}{ (x+1)}\)?

OpenStudy (anonymous):

so the required remainder here should be (-1)^51 +51 =-1+51=50